# On dense orbits in the space of subequivalence relations

### v2

@@ 5 The uniform metric and complexity calculations
Endowing $\operatorname{\mathrm{Sub}}(\mathcal{R})$ with its strong topology and the uniform metric, we get a Polish topometric space in the sense of Ben Yaacov [BY08], and our remark preceding Corollary 4.12 can be upgraded to the fact that if $\mathcal{R}$ has property (T), then $\mathcal{R}$ is a metrically isolated point of $\operatorname{\mathrm{Sub}}(\mathcal{R})$ (see also [GT16] for more examples). I don’t know whether the action of the full group or automorphism group of the hyperfinite ergodic p.m.p. equivalence relation admits a metrically generic orbit, as characterized in [BYM23].

### v3

@@ 5 The uniform metric and complexity calculations
Endowing $\operatorname{\mathrm{Sub}}(\mathcal{R})$ with its strong topology and the uniform metric, we get a Polish topometric space in the sense of Ben Yaacov [BY08], and our remark preceding Corollary 4.12 can be upgraded to the fact that if a p.m.p. equivalence relation $\mathcal{R}$ has property (T), then $\mathcal{R}$ is a metrically isolated point of $\operatorname{\mathrm{Sub}}(\mathcal{R})$ (see also [GT16] for more examples). I don’t know whether the action of the full group or automorphism group of the hyperfinite ergodic p.m.p. equivalence relation admits a metrically generic orbit, as characterized in [BYM23].